Obtain the general solution of the following differential equations.

1. 𝑑𝑦 + 𝑦𝑒x𝑑π‘₯ = 𝑒x𝑑π‘₯

2. π‘₯𝑑𝑦 = (𝑒x βˆ’ 2𝑦)𝑑π‘₯

Obtain the particular solution satisfying the indicated conditions.

3. (2𝑦 βˆ’ π‘₯3)𝑑π‘₯ = π‘₯𝑑𝑦, 𝑦(2) = 4


1
Expert's answer
2021-10-06T10:16:21-0400

1.


"dy=e^x(1-y)dx"

"\dfrac{dy}{1-y}=e^xdx"

Integrate


"\int \dfrac{dy}{1-y}=\int e^xdx"

"-\ln|1-y|=e^x-\ln C"

"1-y=Ce^{-e^x}"

"y=1-Ce^{-e^x}"

2.


"\ud835\udc65\ud835\udc51\ud835\udc66 = (\ud835\udc52^x \u2212 2\ud835\udc66)\ud835\udc51\ud835\udc65"

"y'+2\dfrac{y}{x}=\dfrac{e^x}{x}"

Integrating factor


"\mu(x)=x^2"

"x^2y'+2xy=xe^x"

"d(x^2y)=xe^xdx"

Integrate


"\int d(x^2y)=\int xe^xdx"

"\int xe^xdx"

"\int udv=uv-\int vdu"

"u=x, du=dx"

"dv=e^xdx, v=\int e^x dx=e^x"

"\int xe^xdx=xe^x-\int e^xdx=xe^x-e^x+C"

Then


"x^2y=xe^x-e^x+C"

"y=\dfrac{e^x}{x}-\dfrac{e^x}{x^2}+\dfrac{C}{x^2}"

3.


"y'-\dfrac{2}{x}y=-x^2"

Integrating factor


"\mu(x)=\dfrac{1}{x^2}"

"\dfrac{1}{x^2}y'-\dfrac{2}{x^3}y=-1"

"d(\dfrac{y}{x^2})=-dx"

Integrate


"\int d(\dfrac{y}{x^2})=-\int dx"

"\dfrac{y}{x^2}=-x+C"

"y=-x^3+Cx^2"

"\ud835\udc66(2) = 4"


"4=-(2)^3+C(2)^2=>C=3"

The particular solution satisfying the indicated conditions is

"y=-x^3+3x^2"